OpenStudy (anonymous):

Substitution 6x+2y=0 4x+y=-1

OpenStudy (anonymous):

do you need the process or the answer or both?

OpenStudy (anonymous):

both

OpenStudy (texaschic101):

4x + y = -1 y = -4x - 1 now sub -4x - 1 in for y in the other equation 6x + 2y = 0 6x + 2(-4x - 1) = 0 6x - 8x - 2 = 0 6x - 8x = 2 -2x = 2 x = -1 now sub -1 in for x in either of the equations to find y 4x + y = -1 4(-1) + y = -1 -4 + y = -1 y = -1 + 4 y = 3 check.. 6x + 2y = 0 6(-1) + 2(3) = 0 -6 + 6 = 0 0 = 0 (correct) solution is : (-1,3)

OpenStudy (anonymous):

Thank you @texaschic101

OpenStudy (texaschic101):

no problem...do you have any questions ?